Identify the appropriate distribution function if we want to select 20 random numbers between 0 and 9.


1
Expert's answer
2022-01-31T16:15:44-0500

Solution:

Uniform Distribution is appropriate distribution.

The random-number generator on calculators randomly generates a number between 0 and 9. The random variable X, the number generated, follows a uniform probability distribution.

The notation for the uniform distribution is "X{\sim} U(a, b)" where a= the lowest value of x and b= the highest value of x.

The probability density function is

"f(x)=\frac{1}{b-a}"

for "a \leq x \leq b."

Formulas for the theoretical mean and standard deviation are

"\mu=\dfrac{a+b}2,\sigma=\sqrt{\dfrac{(b-a)^2}{12}}"